OpenStudy (anonymous):

how do you find the polynomial function that has given zeros and a certain condition. EX: -3, 1, and 4; f(2)=30

OpenStudy (jdoe0001):

\(\bf x=-3\implies x+3=0\implies {\color{blue}{ (x+3)}}=0 \\ \quad \\ x=1\implies x-1=0\implies {\color{blue}{ (x-1)}}=0 \\ \quad \\ x=4\implies x-4=0\implies {\color{blue}{ (x-4)}}=0 \\ \quad \\ {\color{red}{ a}}{\color{blue}{ (x+3)}}{\color{blue}{ (x-1)}}{\color{blue}{ (x-4)}}=0\\ \quad \\\implies {\color{red}{ a}}{\color{blue}{ (x+3)}}{\color{blue}{ (x-1)}}{\color{blue}{ (x-4)}}=\textit{original polynomial}\)

OpenStudy (jdoe0001):

so the function will be say like \(\bf y={\color{red}{ a}}{\color{blue}{ (x+3)}}{\color{blue}{ (x-1)}}{\color{blue}{ (x-4)}} \\ \quad \\ f(2)=30\implies x=2\qquad y=30\qquad thus \\ \quad \\ 30={\color{red}{ a}}{\color{blue}{ (2+3)}}{\color{blue}{ (2-1)}}{\color{blue}{ (2-4)}}\) solve for "a" once you get "a", plug it in the expanded equation